OpenStudy (anonymous):

Which numbers are greater than 7 × 10^−4? Choose all answers that are correct. A. 8 × 10^−4 B. 9 × 10^−5 C. 4 × 10^−4 D. 3 × 10^−2 E. 7 × 10^−5

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

A D E?

OpenStudy (anonymous):

A and D.

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@chosenmatt @Kitten_is_back Please help.

OpenStudy (anonymous):

@duck. plz explain

OpenStudy (texaschic101):

rule for negative exponents : a^-4 = 1/a^4 7 * 10^-4 = 7 * 1/10^4 = 7/10000 or 0.0007 A. 8 * 10^-4 = 8 * 1/10^4 = 8/10000 or 0.0008 B. 9 * 10^-5 = 9 * 1/10^5 = 9/100000 or 0.00009 C. 4 * 10^-4 = 4 * 1/10^4 = 4/10000 or 0.0004 D. 7 * 10^-5 = 7 * 1/10^5 = 7/100000 or 0.00007 greater then 0.0007 --- I think it is just A